OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A passes a station at 2:10pm. If train B passes the same station at 2:25pm, what time will train B catch up to train A?

OpenStudy (watchmath):

You forgot to put information about the speed of the two trains.

OpenStudy (anonymous):

oops, train a is traveling at 100 mpr and train B at 120 mph.

OpenStudy (watchmath):

Let \(x\) be the time needed for the traind B to catch up the train A after 2:25 Pm. As train B leaving the station, train A already left the station for 15 minutes =1/4 hour. It means train A is already 1/4*(100)=25 miles from the station. Now the two distance after x hour : distance A = 25 + 100x distance B = 120x the two distance are equal when train B catch up with train A 25+100x=120x 25=20x x=25/20 x=1.25 hour = 1 hour 15 minutes So train B catch up train A at 2:25 + 1 hour 15 minutes = 3:50 pm

OpenStudy (watchmath):

Sorry it should be 3:40 pm

OpenStudy (anonymous):

Thank you!